SCHEDULE: MINI COURSE-2007

Monday  Tuesday  Wednesday  Thursday  Friday
10h-11h: Tignol I  10h-11h: Balmer II  9h-10h: Tignol III  10h-11h: Balmer IV  9h-10h: Vishik V
11h-11h15: coffee  11h-11h15: coffee  10h-10h30: coffee  11h-11h15: coffee  10h-10h30: coffee
11h15-12h15: Tignol II 11h15-12H15: Balmer III  10h30-11H30: Vishik II  11h15-12h15: Tignol IV 10h30-11h30: Balmer V
12h30-14h: lunch  12h30-14h: lunch  11h30-11h45: coffee  12h30-14h: lunch  11h30-11h45: coffee
15h-16h: Balmer I  15h-16h: Vishik I  11h45-12h45: Vishik III  15h-16h: Vishik IV  11h45-12h45: Tignol V
16h-16h30: coffee  16h-16h30: coffee  13h-14h: lunch  16h-16h30: coffee  13h-14h: lunch
16h30-17h30: Pfister  16h30-17h30: Calmès  The Wednesday afternoon is free! 16h30-17h30: Karpenko